# Glassy word problems: ultraslow relaxation, Hilbert space jamming, and computational complexity

### 07 Dec 2023 v1

@@ I Introduction
A common paradigm in quantum dynamics is that isolated quantum systems usually thermalize if one waits long enough Deutsch (1991); Srednicki (1994); Rigol et al. (2008); Nandkishore and Huse (2015); D’Alessio et al. (2016). Indeed, assuming that interactions are spatially local, quantum systems tend to approach a form of local equilibrium on a timescale that is independent of system size, with the late-time dynamics governed by the hydrodynamic relaxation of a small number of conserved densities. The main possible exception to this rule is the many-body localized phase in strongly disordered systems Alet and Laflorencie (2018); Abanin et al. (2019), or in systems which effectively self-generate strong disorder  Schiulaz and Müller (2014); Grover and Fisher (2014); Yao et al. (2016). The structures that can be rigorously shown to arrest thermalization in translation-invariant quantum systems—such as integrability Retore (2022), quantum scars Shiraishi and Mori (2017); Bernien et al. (2017); Turner et al. (2018a, b); Moudgalya et al. (2018); Lin and Motrunich (2019); Schecter and Iadecola (2019); Moudgalya et al. (2022a); Chandran et al. (2023), dynamical constraints Sala et al. (2020a); Khemani et al. (2020a); Moudgalya and Motrunich (2022); Pancotti et al. (2020); van Horssen et al. (2015); Brighi et al. (2023); Lan et al. (2018); Garrahan (2018); De Tomasi et al. (2019); Moudgalya et al. (2022b), etc.—are usually fine-tuned in some sense. However, they are still of experimental relevance since many realistic systems are near the fine-tuned points where these phenomena occur Kohlert et al. (2023); Kinoshita et al. (2006); Scheie et al. (2021); Malvania et al. (2021); Wei et al. (2022); Scherg et al. (2021). Systems near these points have long relaxation timescales and approximate conservation laws that are essentially exact on the timescale of realistic experiments on noisy quantum hardware and cold atom systems. Although the algebraic structure of integrable systems and systems with many-body scars has been well studied, a general understanding of the extent to which local Hilbert space constraints can arrest thermalization is still under development. In this work we introduce an alternative viewpoint for understanding thermalization in a large class of one-dimensional models with local Hilbert space constraints. We begin by developing a general framework for characterizing models with constrained dynamics in terms of semigroups, algebraic structures that resemble groups but need not have inverses or an identity. This approach reproduces examples of constrained models known in the literature, but also provides us with new examples with unusual properties. In particular, it enables us to leverage ideas from the field of geometric group theory to construct (a) models with both exponentially long relaxation times and sharp thermalization transitions, and (b) models where thermalization never occurs, due to an unusual type of ergodicity breaking we call “Hilbert-space jamming” or “fragile fragmentation.” The relation between constrained one-dimensional spin chains and semigroups can be summarized as follows (see Sec. II for a more formal discussion). Most examples of constrained systems in the literature—and all of the examples considered in this work—have constraints which can be formulated in a local product state basis. In these systems, the constraints place certain rules on the processes that the dynamics is allowed to implement, and it is these rules which endow the dynamics with the structure of a semigroup. This identification works by viewing each basis state of the onsite Hilbert space, in the computational basis, as a generator of a semigroup. Since any element of the semigroup can be written as a product of generators, a many-body computational-basis product state is naturally associated with an element of the semigroup, obtained by taking the product of generators from left to right along the chain. We call each computational basis state a word, with each word being a presentation of a certain element in the semigroup. In this picture, the constraints are encoded by requiring that the dynamics preserve the semigroup element associated with each product state. In Sec. II, we show that all local dynamical constraints can be formulated in this way. Of course, not all words represent distinct semigroup elements. For example, in the case where the semigroup is a group $G$ with identity element $\mathtt{e}$ and elements $\mathtt{g}_{1},\mathtt{g}_{2}$, three distinct words of length $4$ representing the same group element are $\mathtt{e}\mathtt{e}\mathtt{e}\mathtt{e}$, $\mathtt{g}_{1}\mathtt{g}_{1}^{-1}\mathtt{g}_{2}^{-1}\mathtt{g}_{2}$, and $\mathtt{g}_{1}\mathtt{g}_{2}\mathtt{g}_{2}^{-1}\mathtt{g}_{1}^{-1}$. Each distinct semigroup element is thus an equivalence class of words under the application of equivalence relations like $\mathtt{g}_{i}\mathtt{g}_{i}^{-1}=\mathtt{e}\mathtt{e}$. The most general local dynamics that preserves semigroup elements is precisely one which locally implements these equivalence relations. The equivalence classes so defined produce multiple sectors of the Hilbert space (“Krylov sectors” or “fragments”) that the dynamical rules are unable to connect, breaking ergodicity and leading to Hilbert space fragmentation Sala et al. (2020a); Khemani et al. (2020a); Moudgalya and Motrunich (2022). Within a given fragment, thermalization of an initial basis state is a process by which the dynamics of the system “works out” which words represent the same semigroup element as the initial state. Crucially, when the problem of determining which words represent a given group element is computationally hard, thermalization within each fragment is slow. This general perspective is powerful because it maps the problem of thermalization in these models onto a well-known algorithmic problem, the word problem for semigroups (a perspective also adopted by Hastings and Freedman in Ref. Hastings and Freedman (2013) to provide examples that exhibit ‘topological obstructions’ and provide a separation between the performances of QMC and quantum annealing). The word problem is the problem referenced above, namely that of deciding whether two words represent the same semigroup element. This identification allows us to lift examples of computationally hard word problems from the mathematical literature to construct models with anomalously slow dynamics. In these models, the dynamics connects the basis states within each fragment in a manner which is much sparser than in generic systems (see Fig. 1 for an illustration), and it is this phenomenon which leads to long thermalization times. The first part of this work focuses on models where the word problem takes exponentially long time to solve. We place particular focus on the “Baumslag-Solitar model”, a spin-2 model with three-site interactions for which the relaxation time of a large class of initial states under any type of dynamics (Hamiltonan, random unitary, etc.) is provably exponentially long in the system size. This model has a conserved charge, and this exponentially long timescale shows up as an exponentially slow hydrodynamic relaxation of density gradients. Not only the relaxation timescale but also the functional form of relaxation is anomalous: a state with density gradients relaxes “gradually, then suddenly,” with an initially prepared density wave experiencing almost no relaxation for exponentially long times, but then undergoing a sudden collapse at a sharply defined onset timescale in a manner reminiscent of the cutoff phenomenon in Markov chains Aldous and Diaconis (1986); Diaconis (1996). Despite this extremely slow hydrodynamics, the states we consider are not dynamically frozen: each configuration is rapidly locally fluctuating, and generic local autocorrelation functions decay rapidly. In the second part of this work, we turn to word problems that have high spatial complexity: i.e., solving them requires not just many steps, but also a large amount of additional space for scratchwork. Put differently, in these problems, deciding whether two words of length $L$ are equivalent requires a derivation involving intermediate words much longer than $L$. In the corresponding dynamical systems, one has (for any fixed $L$) pairs of basis states |$\ket{a},\ket{b}$ such that: (i) |$\ket{a}$ is not connected to |$\ket{b}$ by the dynamics of a chain of length $L$, but (ii) |$\ket{a}\otimes\ket{c}$ is connected to |$\ket{b}\otimes\ket{c}$ by the dynamics of a longer chain, with $|c\rangle$ a fixed ancillary product state. This phenomenon can be viewed in two complementary ways: as a “fragile” form of Hilbert space fragmentation, [^1] In analogy with fragile topology Po et al. (2018), as the ergodicity of the dynamics is modified by the addition of trivial ancillae. or as an unusual type of jamming which has no counterpart in known examples of jammed systems. Using constructions similar to the group model discussed above, we can construct examples where the amount of additional spatial resources grows extremely rapidly with $L$; not just exponentially but also as $e^{e^{L}},e^{e^{e^{L}}}$, and so on. This paper is organized as follows. In Sec. II we introduce word problems for semigroups and groups, and relate them to fragmentation. In Sec. III we use the complexity of the word problem to derive bounds on thermalization times, and in Sec. IV we explore an explicit example of a high-complexity group word problem which yields dynamics with exponentially slow relaxation. We present numerical evidence and analytical estimates for the anomalously slow hydrodynamics of this model. In Sec. V we introduce and analyze a family of group models with fragile fragmentation. Secs. VI and VII respectively present examples based on general semigroups, and generalize our one-dimensional examples to two-dimensional loop models. Finally, we conclude with a discussion of future directions in Sec. VIII.
@@
Figure 1: A schematic illustration of the difference between the Hilbert space connectivity of generic dynamics and group dynamics, with each yellow dot representing a computational basis product state in a single connected sector of the dynamics. In a typical 1D system of length $L$, at most $O(L)$ steps of the dynamics (applications of a Hamiltonian or layers of a unitary circuit) are needed to move between any two basis states (left). In semigroup dynamics (right), the number of steps needed scales as the Dehn function ${\sf Dehn}(L)$ of the semigroup in question, which measures the word problem’s temporal complexity. When ${\sf Dehn}(L)$ is large, the basis states in each sector are connected very sparsely, leading to long thermalization times.
@@ II Semigroup dynamics and constrained 1D systems
In this section we introduce the general framework used to construct the models described above. We will refer to this framework as group dynamics, which encompasses a general class of constrained dynamical systems whose constraints can be derived from the presentation of underlying group or semigroup (to be defined below). These types of constraints are particularly appealing from a theoretical point of view: it turns out that we can rigorously characterize many properties—thermalization times, fragmentation, and so on—using tools from the field of geometric group theory. Broadly speaking, geometric group theory is concerned with characterizing the complexity and geometry of discrete groups (see Clay and Margalit (2017) for an accessible introduction), ideas which will be made precise in the following. As a starting point, we will describe the necessary mathematical background needed to motivate group dynamics. This discussion will center around the word problem, a century-old problem lying at a heart of results regarding the geometry and complexity of groups. We will then see how algorithms solving the word problem can naturally be encoded into the dynamics of 1D spin chains, whose thermalization dynamics is controlled by the word problem’s complexity. Finally, we will see how the structure of the word problem leads to Hilbert space fragmentation, and discuss the properties of the group which control the severity of the fragmentation. Throughout this paper, we will mostly be studying constrained dynamics on 1D spin chains whose onsite Hilbert space is finite dimensional. [^2] Generalizations to 2D are briefly discussed in Sec. VII. We will only consider systems whose time evolution has constraints that can be specified in a local tensor product basis (referred to throughout as the computational basis), either directly or after the application of finite-depth local unitary circuit. In the latter case we will assume the unitary transformation has been done, to avoid loss of generality. We will let $S$ denote the set of computational basis state labels for the onsite Hilbert space, with individual basis states being written as letters in typewriter font ($|\mathtt{a}\rangle,|\mathtt{b}\rangle$, etc.):
@@ II.1 Dynamical constraints and semigroups
Having a dynamical constraint means that not all computational basis product states $|w\rangle,|w^{\prime}\rangle$ can be connected under the dynamics. We will use $\varphi(w)$ to denote the dynamical sector of the state $|w\rangle$, defined as the set of all computational basis product states that $|w\rangle$ can evolve to. We will write ${\sf Dyn}(t)$ to denote time evolution for time $t$ under the dynamics in question, which may be a set of unitary gates, a (possibly space- or time-dependent) Hamiltonian, or a bistochastic Markov chain satisfying the constraints. The existence of the constraint thus means that
$\langle w|{\sf Dyn}(t)|w^{\prime}\rangle\propto\delta_{\varphi(w),\varphi(w^{\prime})}\,\,\forall\,t.$ (2)
@@
Since the tensor product is associative, so too is $\odot$. This equips the set of dynamical sectors with an associative binary operation, thereby endowing it with the structure of a semigroup (a generalization of a group which needn’t have inverses or an identity element). Since all $|w\rangle$ are formed as tensor products of the single-site basis states $|\mathtt{a}\rangle$, the $\varphi(|\mathtt{a}\rangle)$—which we will write simply as $\mathtt{a}$ to save space—constitute a generating set for this semigroup. The dynamical sector of a state $|w\rangle=|\mathtt{a}_{1}\rangle\otimes\cdots\otimes|\mathtt{a}_{L}\rangle$ is then determined simply by multiplying the $\mathtt{a}_{i}$ along the length of the chain [^3] We will mostly be interested in chains with open boundary conditions; for PBC one may arbitrarily choose a site to serve as the “start” of the word.:
@@
Following usual group theory notation, we will denote this semigroup as
@@
where $R$ denotes a set of relations imposed on the product states that can be formed from elements of the on-site computational basis states $S$. The relations in $R$ are determined by $\varphi$, viz. by which product states are related to one another under ${\sf Dyn}$. Consider any two states $|w\rangle$, $|w^{\prime}\rangle$ such that $\varphi(|w\rangle)=\varphi(|w^{\prime}\rangle)$ define the same element of $G$. Since we are interested in dynamics which are geometrically local, it must be possible to relate $|w\rangle$ to $|w^{\prime}\rangle$ using a series of local updates to $|w\rangle$. This means that the set $R$ must be expressible in terms of a set of equivalence relations which each involve only an $O(1)$ number of the elements of $S$—implying in particular that $|R|$ must be finite. A semigroup where both $S$ and $R$ are finite is said to be finitely presented, and all of the semigroups we consider will be of this type. Given a semigroup $G$, we will use the notation ${\sf Dyn}_{G}$ to represent a general local dynamical process acting on $\mathcal{H}=\mathcal{H}_{\rm loc}^{\otimes L}$ which satisfies (2), and hence preserves the dynamical sectors of all computational basis states. The locality of the dynamics means that ${\sf Dyn}_{G}$ must be composed of elementary blocks (unitary gates or Hamiltonian terms) of constant length which, when acting on computational basis product states, implement the relations contained in $R$. Writing the relations in $R$ as $r_{l}=r_{r}$ with $r_{l/r}=\mathtt{a}_{l/r,1}\cdots\mathtt{a}_{l/r,n}$, the locality of ${\sf Dyn}_{G}$ is determined by the maximal length of the $r_{l/r}$, which we denote as $\ell_{R}$:
@@
where $|r_{l/r}\rangle_{i}=|\mathtt{a}_{l/r,1}\rangle_{i}\otimes\cdots\otimes|\mathtt{a}_{l/r,n}\rangle_{i+n}$ and the $\lambda_{i,r}$ are arbitrary complex numbers. Note that the above Hamiltonian is only well-defined if $|r_{l}|=|r_{r}|$ for all relations $r$. In cases where this does not hold, we will rectify this problem by adding a trivial character e to the onsite Hilbert space—with $\mathtt{e}$ defined to commute with all of the other generators $\mathtt{a}$—which allows us to then ‘pad’ the relations $r$ in a way which ensures that $|r_{l}|=|r_{r}|$. As a concrete example, consider the semigroup $\mathbb{Z}^{2}=\langle\mathtt{x},\mathtt{y}|\,\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}\rangle$, [^4] When writing presentations of groups (such as $\mathbb{Z}^{2}$), we will omit the inverse generators (here $\mathtt{x}^{-1},\mathtt{y}^{-1}$) and the identity $\mathtt{e}$ to save space. In this example, the full generating set is $S=\{\mathtt{x},\mathtt{y},\mathtt{x}^{-1},\mathtt{y}^{-1},\mathtt{e}\}$. which as we will see later in some sense has trivial dynamics. Since the generating set of this presentation has dimension $|S|=5$, a Hamiltonian $H_{\mathbb{Z}^{2}}$ with $\mathbb{Z}^{2}$-constrained dynamics thus acts most naturally on a spin-2 chain. The single nontrivial relation $\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}$ has length $\ell_{R}=2$, and thus $H_{\mathbb{Z}^{2}}$ can be taken to be 2-local, assuming a form like
@@
Figure 2: Semigroup dynamics and the word problem. a) In the word problem, we are given a length-$L$ word $|w\rangle$ and a series of re-writing rules that let us make local updates to the characters of the word. The word problem for $|w\rangle$ is the task of deciding whether or not a sequence of allowed updates can be found which transforms $|w\rangle$ into a particular reference word, here chosen to be $|\mathtt{e}^{L}\rangle$. The Dehn function ${\sf Dehn}(w)$ measures the time complexity of the word problem, viz. the minimal number of updates needed to connect $|w\rangle$ to $|\mathtt{e}^{L}\rangle$. b) In some situations, $|w\rangle$ can only be transformed into $|\mathtt{e}^{L}\rangle$ by increasing the amount of available space, done here by appending $|\mathtt{e}\mathtt{e}\cdots\rangle$ onto the original word. The minimal amount of extra space needed defines the expansion length ${\sf EL}(w)$, which captures the spatial complexity of the word problem. c) Our group dynamics models are defined on a 1D chain, with an onsite Hilbert space has a basis we label as $\mathcal{H}_{\rm loc}=\{|\mathtt{g}\rangle,|\mathtt{g}^{-1}\rangle,|\mathtt{e}\rangle\}$; here $\mathtt{g}\in S$ run over the generators of a discrete group $G$, and $\mathtt{e}$ represents the identity element. The dynamics in these models is restricted to local updates which preserve the group element obtained by multiplying the generators along the length of the chain.
@@
The construction of group-constrained random unitary dynamics is similar to the Hamiltonian case. For random unitary dynamics, ${\sf Dyn}_{G}$ is constructed using $\ell_{R}$-site unitary gates $U$ whose matrix elements $\langle w^{\prime}|U|w\rangle$ are nonzero only if the length-$\ell_{R}$ words $w,w^{\prime}$ satisfy $\varphi(w)=\varphi(w^{\prime})$. Such unitaries admit the decomposition
$U=\bigoplus_{g\in G_{\ell_{R}}}U_{g},\hskip 14.22636ptG_{\ell}=\{g\,|\,\exists\,w\,:\,\varphi(w)=g,\,|w|\leq\ell_{R}\},$ (9)

### 11 Apr 2024 v2

@@ I Introduction
A common paradigm in quantum dynamics is that isolated quantum systems usually thermalize if one waits long enough [1, 2, 3, 4, 5]. Indeed, assuming that interactions are spatially local, quantum systems tend to approach a form of local equilibrium on a timescale that is independent of system size, with the late-time dynamics governed by the hydrodynamic relaxation of a small number of conserved densities. The main possible exception to this rule is the many-body localized phase in strongly disordered systems [6, 7], or in systems which effectively self-generate strong disorder  [8, 9, 10]. The structures that can be rigorously shown to arrest thermalization in translation-invariant quantum systems—such as integrability [11], quantum scars [12, 13, 14, 15, 16, 17, 18, 19, 20], dynamical constraints [21, 22, 23, 24, 25, 26, 27, 28, 29, 30], etc.—are usually fine-tuned in some sense. However, they are still of experimental relevance since many realistic systems are near the fine-tuned points where these phenomena occur [31, 32, 33, 34, 35, 36]. Systems near these points have long relaxation timescales and approximate conservation laws that are essentially exact on the timescale of realistic experiments on noisy quantum hardware and cold atom systems. Although the algebraic structure of integrable systems and systems with many-body scars has been well studied, a general understanding of the extent to which local Hilbert space constraints can arrest thermalization is still under development. In this work we introduce an alternative viewpoint for understanding thermalization in a large class of one-dimensional models with local Hilbert space constraints. We begin by developing a general framework for characterizing models with constrained dynamics in terms of semigroups, algebraic structures that resemble groups but need not have inverses or an identity. This approach reproduces examples of constrained models known in the literature, but also provides us with new examples with unusual properties. In particular, it enables us to leverage ideas from the field of geometric group theory to construct (a) models with both exponentially long relaxation times and sharp thermalization transitions, and (b) models where thermalization never occurs due to an unusual type of ergodicity breaking we call “Hilbert-space jamming” or “fragile fragmentation.” The relation between constrained one-dimensional spin chains and semigroups can be summarized as follows (see Sec. II for a more formal discussion). Most examples of constrained systems in the literature—and all of the examples considered in this work—have constraints which can be formulated in a local product state basis. In these systems, the constraints place certain rules on the processes that the dynamics is allowed to implement, and it is these rules which endow the dynamics with the structure of a semigroup. This identification works by viewing each basis state of the onsite Hilbert space, in the computational basis, as a generator of a semigroup. Since any element of the semigroup can be written as a product of generators, a many-body computational-basis product state is naturally associated with an element of the semigroup, obtained by taking the product of generators from left to right along the chain. We call each computational basis state a word, with each word being a presentation of a certain element in the semigroup. In this picture, the constraints are encoded by requiring that the dynamics preserve the semigroup element associated with each product state. In Sec. II, we show that all local dynamical constraints can be formulated in this way. Of course, not all words represent distinct semigroup elements. For example, in the case where the semigroup is a group $G$ with identity element $\mathtt{e}$ and elements $\mathtt{g}_{1},\mathtt{g}_{2}$, three distinct words of length $4$ representing the same group element are $\mathtt{e}\mathtt{e}\mathtt{e}\mathtt{e}$, $\mathtt{g}_{1}\mathtt{g}_{1}^{-1}\mathtt{g}_{2}^{-1}\mathtt{g}_{2}$, and $\mathtt{g}_{1}\mathtt{g}_{2}\mathtt{g}_{2}^{-1}\mathtt{g}_{1}^{-1}$. Each distinct semigroup element is thus an equivalence class of words under the application of equivalence relations like $\mathtt{g}_{i}\mathtt{g}_{i}^{-1}=\mathtt{e}\mathtt{e}$. The most general local dynamics that preserves semigroup elements is precisely one which locally implements these equivalence relations. The equivalence classes so defined produce multiple sectors of the Hilbert space (“Krylov sectors” or “fragments”) that the dynamical rules are unable to connect, breaking ergodicity and leading to Hilbert space fragmentation [21, 22, 23]. Within a given fragment, thermalization of an initial basis state is a process by which the dynamics of the system “works out” which words represent the same semigroup element as the initial state. Crucially, when the problem of determining which words represent a given element is computationally hard, thermalization within each fragment is slow. This general perspective is powerful because it maps the problem of thermalization in these models onto a well-known algorithmic problem, the word problem for semigroups (a perspective also adopted by Hastings and Freedman in Ref. [37] to provide examples of dynamics exhibiting ‘topological obstructions’ which provide a separation between the performances of QMC and quantum annealing). The word problem is the problem referenced above, namely that of deciding whether two words represent the same semigroup element. This identification allows us to lift examples of computationally hard word problems from the mathematical literature to construct models with anomalously slow dynamics. In these models, the dynamics connects the basis states within each fragment in a manner which is much sparser than in generic systems (see Fig. 1 for an illustration), and it is this phenomenon which leads to long thermalization times. The first part of this work focuses on models where the word problem takes an exponentially long time to solve. We place particular focus on the “Baumslag-Solitar model”, a spin-2 model with three-site interactions for which the relaxation time of a large class of initial states under any type of local dynamics (Hamiltonan, random unitary, etc.) is provably exponentially long in the system size. This model has a conserved charge, and this exponentially long timescale shows up as an exponentially slow hydrodynamic relaxation of density gradients. Not only the relaxation timescale but also the functional form of relaxation is anomalous: a state with density gradients relaxes “gradually, then suddenly,” with an initially prepared density wave experiencing almost no relaxation for exponentially long times, but then undergoing a sudden collapse at a sharply defined timescale. Despite this extremely slow hydrodynamics, the states involved are not dynamically frozen: each configuration is rapidly locally fluctuating, and generic local autocorrelation functions decay rapidly. In the second part of this work, we turn to word problems that have high spatial complexity: i.e., solving them requires not just many steps, but also a large amount of additional space for scratchwork. Put differently, in these problems, deciding whether two words of length $L$ are equivalent requires a derivation involving intermediate words much longer than $L$. In the corresponding dynamical systems, one has (for any fixed $L$) pairs of basis states |$\ket{a},\ket{b}$ such that: (i) |$\ket{a}$ is not connected to |$\ket{b}$ by the dynamics of a chain of length $L$, but (ii) |$\ket{a}\otimes\ket{c}$ is connected to |$\ket{b}\otimes\ket{c}$ by the dynamics of a longer chain, with $|c\rangle$ a fixed ancillary product state. This phenomenon can be viewed in two complementary ways: as a “fragile” form of Hilbert space fragmentation, [^1] In analogy with fragile topology [40], as the ergodicity of the dynamics is modified by the addition of trivial ancillae. or as an unusual type of jamming which has no counterpart in known examples of jammed systems. Using constructions similar to the group model discussed above, we can construct examples where the amount of additional spatial resources grows extremely rapidly with $L$; not just exponentially but also as $e^{e^{L}},e^{e^{e^{L}}}$, and so on. This paper is organized as follows. In Sec. II we introduce word problems for semigroups and groups, and relate them to fragmentation. In Sec. III we use the complexity of the word problem to derive bounds on thermalization times, and in Sec. IV we explore an explicit example of a high-complexity group word problem which yields dynamics with exponentially slow relaxation. We present numerical evidence and analytical estimates for the anomalously slow hydrodynamics of this model. In Sec. V we introduce and analyze a family of group models with fragile fragmentation. Secs. VI and VII respectively present examples based on semigroups, and generalize our one-dimensional examples to two-dimensional loop models. Finally, we conclude with a discussion of future directions in Sec. VIII.
@@
Figure 1: A schematic illustration of the difference between the Hilbert space connectivity of generic dynamics and semigroup dynamics, with each yellow dot representing a computational basis product state in a single connected sector of the dynamics. In a 1D system of length $L$, typical dynamics (left) requires at most $O(L)$ steps of the dynamics (applications of a Hamiltonian or layers of a unitary circuit) to move between any two basis states. In semigroup dynamics (right), the number of steps needed scales as the Dehn function ${\sf Dehn}(L)$ of the semigroup in question, which measures the word problem’s temporal complexity. When ${\sf Dehn}(L)$ is large, the basis states in each sector are connected very sparsely, leading to long thermalization times.
@@ II Semigroup dynamics and constrained 1D systems
In this section we introduce the general framework used to construct the models described above. We will refer to this framework as semigroup dynamics, which encompasses a general class of constrained dynamical systems whose constraints can be derived from the presentation of underlying group or semigroup (to be defined below). These types of constraints are particularly appealing from a theoretical point of view: it turns out that we can rigorously characterize many properties—thermalization times, fragmentation, and so on—using tools from the field of geometric group theory. Broadly speaking, geometric group theory is concerned with characterizing the complexity and geometry of discrete groups (see [41] for an accessible introduction), ideas which will be made precise in the following. As a starting point, we will describe the necessary mathematical background needed to motivate group dynamics. This discussion will center around the word problem, a century-old problem lying at a heart of results regarding the geometry and complexity of groups. We will then see how algorithms solving the word problem can naturally be encoded into the dynamics of 1D spin chains, whose thermalization dynamics is controlled by the word problem’s complexity. Finally, we will see how the structure of the word problem leads to Hilbert space fragmentation, and discuss the properties of the group which control the severity of the fragmentation. Throughout this paper, we will mostly be studying constrained dynamics on 1D spin chains whose onsite Hilbert space is finite dimensional. [^2] Generalizations to 2D are briefly discussed in Sec. VII. We will only consider systems whose time evolution has constraints that can be specified in a local tensor product basis (referred to throughout as the computational basis), either directly or after the application of finite-depth local unitary circuit. In the latter case we will assume the unitary transformation has been done, to avoid loss of generality. We will let $S$ denote the set of computational basis state labels for the onsite Hilbert space, with individual basis states being written as letters in typewriter font ($|\mathtt{a}\rangle,|\mathtt{b}\rangle$, etc.):
@@ II.1 Dynamical constraints and semigroups
We will write ${\sf Dyn}(t)$ to denote time evolution for time $t$ under the dynamics in question, which may be performed using a set of unitary gates, a (possibly space- or time-dependent) Hamiltonian, or a bistochastic Markov chain. Having a dynamical constraint means that not all computational basis product states $|w\rangle,|w^{\prime}\rangle$ can be connected under the dynamics. We will use $\varphi(|w\rangle)$ to denote the dynamical sector of the state $|w\rangle$, defined as the set of all computational basis product states that $|w\rangle$ can evolve to, thus
$\langle w|{\sf Dyn}(t)|w^{\prime}\rangle\propto\delta_{\varphi(|w\rangle),\varphi(|w^{\prime}\rangle)}\,\,\forall\,t.$ (2)
@@
Since the tensor product is associative, so too is $\odot$. This equips the set of dynamical sectors with an associative binary operation, thereby endowing it with the structure of a semigroup, a generalization of a group which needn’t have inverses or an identity element. [^3] We focus on the more general structure of a semigroup (as opposed to a group), since it provides the most general framework for discussing 1D constrained dynamics, does not complicate the bounds we will obtain on thermalization dynamics, and because most of the existing models studied in the literature are described by semigroups, rather than groups. Since all $|w\rangle$ are formed as tensor products of the single-site basis states $|\mathtt{a}\rangle$, the $\varphi(|\mathtt{a}\rangle)$—which we will write simply as $\mathtt{a}$ to save space—constitute a generating set for this semigroup. The dynamical sector of a state $|w\rangle=|\mathtt{a}_{1}\rangle\otimes\cdots\otimes|\mathtt{a}_{L}\rangle$ is then determined simply by multiplying the $\mathtt{a}_{i}$ along the length of the chain [^4] We will mostly be interested in chains with open boundary conditions; for PBC one may arbitrarily choose a site to serve as the “start” of the word.:
@@
Following usual group theory notation, we will denote a semigroup with generating set $S$ as
@@
where $R$ denotes a set of relations imposed on the product states that can be formed from elements of the on-site computational basis states $S$. When writing presentations of groups we will omit the inverse generators and the identity from $S$, and will likewise omit trivial relations like $\mathtt{a}\mathtt{a}^{-1}=\mathtt{e},\mathtt{a}\mathtt{e}=\mathtt{e}\mathtt{a}$ from $R$. For the remainder of the paper, presentations of semigroups which are not groups will always be denoted by ${\sf semi}\langle S\,|R\rangle$, while presentations of groups will be denoted simply by $\langle S\,|\,R\rangle$. The relations in $R$ are determined by $\varphi$, namely by which product states are related to one another under ${\sf Dyn}$. Consider any two states $|w\rangle$, $|w^{\prime}\rangle$ such that $\varphi(|w\rangle)=\varphi(|w^{\prime}\rangle)$ define the same element of $G$. Since we are interested in dynamics which are geometrically local, it must be possible to relate $|w\rangle$ to $|w^{\prime}\rangle$ using a series of local updates to $|w\rangle$. This means that the set $R$ must be expressible in terms of a set of equivalence relations which each involve only an $O(1)$ number of the elements of $S$—implying in particular that $|R|$ must be finite. A semigroup where both $S$ and $R$ are finite is said to be finitely presented, and all of the semigroups we consider will be of this type. Given a semigroup $G$, we will use the notation ${\sf Dyn}_{G}$ to represent a general local dynamical process acting on $\mathcal{H}=\mathcal{H}_{\rm loc}^{\otimes L}$ which satisfies the constraint (2), and hence preserves the dynamical sectors of all computational basis states. The locality of the dynamics means that ${\sf Dyn}_{G}$ must be composed of elementary blocks (unitary gates or Hamiltonian terms) of constant length which, when acting on computational basis product states, implement the relations contained in $R$. Writing the relations in $R$ as $r_{l}=r_{r}$ with $r_{l/r}=\mathtt{a}_{l/r,1}\cdots\mathtt{a}_{l/r,n}$, the locality of ${\sf Dyn}_{G}$ is determined by the maximal length of the $r_{l/r}$, which we denote as $\ell_{R}$:
@@
where $|r_{l/r}\rangle_{i}=|\mathtt{a}_{l/r,1}\rangle_{i}\otimes\cdots\otimes|\mathtt{a}_{l/r,n}\rangle_{i+n}$ and the $\lambda_{i,r}$ are arbitrary complex numbers. Note that the above Hamiltonian is only well-defined if $|r_{l}|=|r_{r}|$ for all relations $r$. In cases where this does not hold, we will rectify this problem by adding a trivial character e to the onsite Hilbert space—with $\mathtt{e}$ defined to commute with all of the other generators $\mathtt{a}$—which allows us to then ‘pad’ the relations $r$ in a way which ensures that $|r_{l}|=|r_{r}|$. As a simple example, consider the group $\mathbb{Z}^{2}=\langle\mathtt{x},\mathtt{y}|\,\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}\rangle$, which as we will see later in some sense has trivial dynamics. Since the full generating set $S=\{\mathtt{x},\mathtt{x}^{-1},\mathtt{y},\mathtt{y}^{-1},\mathtt{e}\}$ of this presentation has dimension $|S|=5$, a Hamiltonian $H_{\mathbb{Z}^{2}}$ with $\mathbb{Z}^{2}$-constrained dynamics thus acts most naturally on a spin-2 chain. The single nontrivial relation $\mathtt{x}\mathtt{y}=\mathtt{y}\mathtt{x}$ has length $\ell_{R}=2$, and thus $H_{\mathbb{Z}^{2}}$ can be taken to be 2-local, assuming a form like
@@
Figure 2: Semigroup dynamics and the word problem. a) In the word problem, we are given a length-$L$ word $|w\rangle$ and a series of re-writing rules that let us make local updates to the characters of the word. The word problem for $|w\rangle$ is the task of deciding whether or not a sequence of allowed updates can be found which transforms $|w\rangle$ into a particular reference word, here chosen to be $|\mathtt{e}^{L}\rangle$. The Dehn function ${\sf Dehn}(w)$ measures the time complexity of the word problem, namely the minimal number of updates needed to connect $|w\rangle$ to $|\mathtt{e}^{L}\rangle$. b) In some situations, $|w\rangle$ can only be transformed into $|\mathtt{e}^{L}\rangle$ by increasing the amount of available space, done here by appending $|\mathtt{e}\mathtt{e}\cdots\rangle$ onto the original word. The minimal amount of extra space needed defines the expansion length ${\sf EL}(w)$, which captures the spatial complexity of the word problem. c) For a semigroup $G=\langle S|R\rangle$ defined by generators $\mathtt{a}\in S$ and relations between the generators $r_{i}\in R$, our construction defines a dynamics ${\sf Dyn}_{G}$ which acts on a 1D chain with an onsite Hilbert space $\mathcal{H}_{\rm loc}=\{|\mathtt{a}\rangle\,:\,\mathtt{a}\in S\}$. The dynamics implimented by ${\sf Dyn}_{G}$ is restricted to local updates which preserve the semigroup element obtained by multiplying the generators along the length of the chain; the allowed updates are consequently fixed by the relations in $R$ (with the figure drawn using the relations $r_{1}:\mathtt{a}\mathtt{b}=\mathtt{c}\mathtt{d}$, $r_{2}:\mathtt{e}\mathtt{f}=\mathtt{g}\mathtt{h}$, etc.).
@@
The construction of group-constrained random unitary dynamics is similar to the Hamiltonian case. For random unitary dynamics, ${\sf Dyn}_{G}$ is constructed using $\ell_{R}$-site unitary gates $U$ whose matrix elements $\langle w^{\prime}|U|w\rangle$ are nonzero only if the length-$\ell_{R}$ words $w,w^{\prime}$ satisfy $\varphi(|w\rangle)=\varphi(|w^{\prime}\rangle)$. Such unitaries admit the decomposition
$U=\bigoplus_{g\in G_{\ell_{R}}}U_{g},\hskip 8.5359ptG_{\ell_{R}}=\{g\,|\,\exists\,|w\rangle\,:\,\varphi(|w\rangle)=g,\,|w|\leq\ell_{R}\},$ (9)